OpenStudy (lukecrayonz):

Find (d^2y)/(dx^2) y=9x-8 Please explain because I just started this:(

OpenStudy (campbell_st):

can you find the 1st derivative..? what is it?

OpenStudy (lukecrayonz):

9

OpenStudy (lukecrayonz):

@campbell_st

OpenStudy (campbell_st):

ok... so you need to differentiate the 1st derivative to get the 2nd derivative.. what happens when you differentiate a constant..?

OpenStudy (lukecrayonz):

I understand to find the deriv of the deriv, but is my final answer going to be 9?

OpenStudy (lukecrayonz):

0 i meant zero for the last word

OpenStudy (campbell_st):

nope... the power rule is \[y = x^n... y' = nx^{n-1}\] so think of it as \[y' = 9x^0\] use the power rule.. \[y'' = 0 \times 9x^{0-1}\] and zero times anything is zero...

OpenStudy (lukecrayonz):

I had a typo :P I meant zero. But thank you. I didn't know it was simply finding the deriv of the deriv. So for fourth deriv, it's just the deriv of the deriv of the deriv of the deriv?

OpenStudy (campbell_st):

thats ok... just remember the derivative of a constant is zero. thats correct... but you would normally only be asked to find the 1st and 2nd derivatives... as they are critical in curve sketching and rates of change

OpenStudy (lukecrayonz):

@campbell_st question :) f(x)=(x^2+8)^8 f''(x)

OpenStudy (campbell_st):

ok... you need the chain rule for the 1st derivative then the 2nd derivative will need the product rule and chain rule... any thoughts on the 1st derivative..

OpenStudy (lukecrayonz):

So I take (x^2+8)^8 g(x)=x^2+8 k=8 8(x^2+8)^7*2x

OpenStudy (campbell_st):

thats correct... tidy it up a bit and you get dy/dx = 16x(x^2 + 8)^7 now you need to product and chain rules to get the 2nd derivative

OpenStudy (lukecrayonz):

Can you help me with that part? I am horrible at product rule.

OpenStudy (lukecrayonz):

Can I just distribute 16x through the parentheses, and get (16x^3+128x)^7 then do chain rule?

OpenStudy (campbell_st):

nope... doesn't work like that... because you have 7 lots of (x^2 + 8) so you would need to do it 7 times and then distribute to get a degree 21 polynomial

OpenStudy (lukecrayonz):

So wait, my f(x) is 16x and my g(x) is (x^2+8)^7?

OpenStudy (lukecrayonz):

I've only done quotient rules of fractions

OpenStudy (campbell_st):

thats correct... the product rule is y' = f(x) * g'(x) + f'(x)*g(x)

OpenStudy (lukecrayonz):

I made a new question, i'll tag you.

OpenStudy (campbell_st):

there are plenty of people who can help... no need to tag me